OpenStudy (abmon98):

((3)^1/2)*tan(theta)=2sin(theta) solve the equation for the domain 0< or equal to (theta)< or equal to 2Pi

OpenStudy (abmon98):

\[\sqrt{3} \tan \theta=2\sin \theta, 0\le \theta \le360 \]

OpenStudy (anonymous):

\[ \sqrt{3} \tan (\theta )=2 \sin (\theta )\\ \frac{\sqrt{3} \sin (\theta )}{\cos (\theta )}-2 \sin (\theta )=0\\ \sin (\theta ) \left(\frac{\sqrt{3}}{\cos (\theta )}-2\right)=0\\ \implies \sin (\theta )=0 \\ or\\ \frac{\sqrt{3}}{\cos (\theta )}=2\\ \cos (\theta )=\frac{\sqrt{3}}{2} \] The rest should be easy

OpenStudy (anonymous):

The above gives rise to \[ \theta = 0\\ \theta = \pi\\ \theta = 2\pi\\ \theta = \frac \pi 6\\ \theta =- \frac \pi 6+ 2 \pi =\frac {11 \pi}6 \]

OpenStudy (abmon98):

why is theta equal to 0 and 2pi

OpenStudy (anonymous):

The sin is equal zero at 0, pi and 2 pi

OpenStudy (anonymous):

The roots I stated above are all between 0 and 2 pi included

OpenStudy (abmon98):

oh ok thank you :D